Would the following solved for the currents in each series, the resistance is to be divided by 100, so on the first one for example the resistances are 5, 7.5, and 10, everything else is the same11 R1 (500) A А A R2 (750 (2) E (10.0 V) 12 13 R3 (1000 (2) €2 (5.00 V)EA (5.00 V) E2 (10.0 V) 12 Re (800 2) 1 R4 (500 ) N A A 위 A 13 A Ri (1000 ) R3 (600 )
Answer #1

Part-a3 را ہے۔ 2 A А 8,(10.00) R (500) V Iz ] R2(7580 & croorge & R3 (1000) Kirchhoffs law Applying Kirchoffs in loop O- lo5- SI -7.5 I 220 I, Izt Iz 5-55/2 513 -7,5 I 2 55, -5113,-1013 30 t 2 5 +2.5 I3=0 ² I 3-5 = 2 A U I 3 equation E from 10х3 7.Iz f E, C51001 R₂(8002) &z(10.04) Ş RESTO ITA Roche) A B/160) in loop 5+ I2 R2 + IIR, = 0 A o 5+ 8 I 2 + 10 II- in loop @ s t15 - 613 + 10 I2 7101q 1 o I 3=0 or 15 +1012 - 1653+1054 =0 F 4 or G It In + 2 Izt 214 - 2I3 =0 1+ 3 I 4 +212 2 I 3 = 0 & I 2from сем G It 3 2 so 13 ) + 2/22-13-2) 18 Iz + 4 +13 It 25 cally 18 7 - 25 + 2212 Iz g أي 63-25 1oI3-162. I 3 4.5 T 38 در) 52direction from -> H 18 Iz E 22 X 3-41 25 al I = 2:78 A 12 A from I, + 3.41 = 2.98 2:09 I - 2.72 A direction is also opposite